OpenStudy (anonymous):

Let f(x) =3x-1 find f[f(x)]

OpenStudy (anonymous):

If f(x) = 3x-1 f[f(x)] you simply replace your x with your function so f(f(x)) = 3(3x-1)-1

OpenStudy (anonymous):

\[f(f(x)) = 3(f(x)) -1 \] \[f(f(x)) = 3(3x-1) - 1\] \[f(f(x))= 9x - 3 -1 = 9x -4\]

OpenStudy (anonymous):

f[f(x)]=9x-3-1= 9x-4

OpenStudy (anonymous):

cool that is what I got as well